Write the equation of the line with given slope and y-intercept. Then graph each the line using the slope and y-intercept.
Answer by fractalier(6550)

Of course you did not supply the specific slope and y-intercept, but if you had, this is what you would do...
Given m = 2/3 and y-int = -4
you plug into the slope-intercept form
y = mx + b
where 2/3 goes in for m and -4 goes in for b...giving you
y = (2/3)x - 4
To graph it, put a point on the y-axis at -4...the actual point is (0, -4)...from there, then, go up two and to the right three (a slope of 2/3) and put another point...then connect 'em and you're done...